Mcdale Inc. produces and sells two products. Data concerning those products for the most recent month appear below:

Product I49V Product Z50U
Sales $ 50,000 $ 55,000
Variable expenses $ 13,800 $ 27,150

The fixed expenses of the entire company were $39,040. The break-even point for the entire company is closest to:
Answer #1

Contribution margin = Total sales - Total Variable cost

= (50,000+55,000) - (13,800+27,150)

= 105,000 - 40,950

= 64,050

Contribution margin ratio = 64,050/105,000 = 61%

Breakeven point = Fixed cost/Contribution margin ratio

= 39,040/61%

= 64,000